OpenStudy (darkbeautystar):

Find the Area of a regular octagon with an apothem of 8 and a side length of 8.93.

OpenStudy (darkbeautystar):

@TheSmartOne

OpenStudy (darkbeautystar):

@dan815

OpenStudy (anonymous):

All you really need is the side length. A=2(1+sqrt(2))x^2 Fill 8.93 in for x and you get an area of 385.04.

Directrix (directrix):

The area of a regular polygon can be found using the formula: A = (1/2) * a * p where a is the apothem and p is the perimeter of the polygon. An octagon has 8 sides. Because the polygon is regular, the polygon is equilateral and all 8 sides have the same length.

Directrix (directrix):

a = 8 p = 8 * 8.93 A =(1/2) * a * p A = (1/2) * 8 * ( 8 * 8.93) A = 285.76 @darkbeautystar